# Entropy Functions on Two-Dimensional Faces of Polymatroidal Region of Degree Four: Part I: Problem Formulation and More

### v3

@@ I Introduction
By definition, $\Gamma_{n}$ is a polyhedral cone in $\mathcal{H}_{n}$. Thus, each Shannon-type information inequality determines a face $F$ of $\Gamma_{n}$. It is natural to characterize entropy functions on the specific $F$ of $\Gamma_{n}$ (See Subsection II-B for details on the faces of a polyhedral cone). Let $F^{*}\triangleq F\cap\Gamma^{*}_{n}$ be the set of all entropy functions in $F$. In what follows, when we refer to determining the entropy functions on $F$ , or the region $F^{*}$, we will simply call this characterizing $F$. A non-Shannon-type information inequality can be considered as an outer bound on $F^{*}$ when $F=\Gamma_{n}$ itself, the improper face of $\Gamma_{n}$. A constrained non-Shannon-type information inequality is an outer bound on $F^{*}$ when $F$ is the face determined by the constraints that are equalities obtained by setting some Shannon-type inequalities as equalities. When $F$ is an extreme ray, i.e., a $1$-dimensional face of $\Gamma_{n}$, if it contains a matroid, entropy functions on $F$ are called matroidal entropy functions, and they can be fully characterized by the probabilistically characteristic set of the matroid [7, 8, 9]. Matúš fully characterized the first non-trivial 2-dimensional face of $\Gamma_{n}$ in 2006 [10]. It is a $2$-dimensional face of $\Gamma_{3}$. In 2012, Chen and Yeung characterized another $2$-dimensional face of $\Gamma_{3}$ [11]. They are the only two types of non-trivial $2$-dimensional faces of $\Gamma_{3}$ that need to be characterized. To the best of our knowledge, so far, there is no fully characterized no-trivial $3$-dimensional faces. However, partial characterizations of $3$-dimensional faces of $\Gamma_{3}$ can be found in [12, 13].
@@ II-B Faces of a polyhedral cone
$\mathbf{r}(A)=\begin{cases}\min\{2,|A|\}\quad&A\neq\{i\}\\ 2\quad&A=\{i\}\end{cases}$
@@ II-C Entropy functions on the extreme rays of $\Gamma_{4}$
$U^{\alpha}_{2,3}$ II-C.2 [Uncaptioned image] $\{a\mathbf{r}:a=\log v,\ v>0,v\in\mathbb{Z}\}$
$\mathcal{W}^{\alpha}_{2}$ II-C.2 [Uncaptioned image] $\{a\mathbf{r}:a=\log v,\ v>0,v\in\mathbb{Z}\}$
$U_{2,4}$ II-C.3 [Uncaptioned image] $\{a\mathbf{r}:a=\log v,v>0,v\in\mathbb{Z},v\neq 2,6\}$
$U_{3,4}$ II-C.2 [Uncaptioned image] $\{a\mathbf{r}:a=\log v,\ v>0,v\in\mathbb{Z}\}$
$\hat{U}^{i}_{2,5}$ II-C.4 [Uncaptioned image] $\{a\mathbf{r}:a=\log v,\ v>0,v\in\mathbb{Z}\}$
$\hat{U}^{i}_{3,5}$ II-C.5 [Uncaptioned image] $\{a\mathbf{r}:a=\log v,\ v>0,v\in\mathbb{Z}\}$
@@
For the rank function $\mathbf{r}$ of $\hat{U}^{i}_{2,5}$, $\mathbf{h}=a\mathbf{r}$ is entropic if and only if $a=\log v$ for integer $v>0$.
@@
For $\mathbf{h}=a\mathbf{r}$, where $\mathbf{r}$ is the rank function of $\hat{U}^{1}_{2,5}$, if we restrict it on $\{2,3,4\}$, we will see that its restriction $\mathbf{h}^{\prime}=a\mathbf{r}_{1}$, where $\mathbf{r}_{1}$ is the rank function of $U_{2,3}$, is entropic only if $a=\log v$ for integer $v>0$.
@@
For the rank function $\mathbf{r}$ of $\hat{U}^{i}_{3,5}$, $\mathbf{h}=a\mathbf{r}$ is entropic if and only if $a=\log v$ for integer $v>0$.
@@
By (II.C.24), (II.C.41), (II.C.34), (II.C.35), (II.C.43), (II.C.39), (II.C.40) and (II.C.42) the value of $H(X_{A})$ are determined for all $A\subseteq N$ except for $A=\{1\}$. Now we verify that $H(X_{1})=2\log v$. By (II.C.24) and (II.C.41),
@@ IV-B Entropy functions on the faces extended from $2$-dimensional faces of $\Gamma_{3}$
For $F=(U_{2,3}^{123},U_{1,1}^{1})$, $(U_{2,3}^{123},U_{1,1}^{4})$, $(U_{2,3}^{123},U_{1,2}^{14})$, $(\mathcal{W}_{2}^{14},U_{1,1}^{1})$, $(\mathcal{W}_{2}^{34},U_{1,1}^{1})$, $(\mathcal{W}_{2}^{14},U_{1,2}^{14})$ and $(\mathcal{W}_{2}^{24},U_{1,2}^{14})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a=\log{v}$ for integer $v>0$.
@@
For $F=(\hat{U}_{2,5}^{1},U_{1,1}^{1})$, $(\hat{U}_{2,5}^{1},U_{1,1}^{2})$ and $(\hat{U}_{2,5}^{1},U_{1,2}^{12})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a=\log v$ for integer $v>0$.
@@
By restriction, the above $3$ faces of $\Gamma_{4}$ are mapped to a subset of $\Gamma_{3}$ on $\{2,3,4\}$. The face $(\hat{U}_{2,5}^{1},U_{1,1}^{1})$ of $\Gamma_{4}$ is mapped to the extreme ray $U_{2,3}$ of $\Gamma_{3}$ on $\{2,3,4\}$, which implies that $a$ can only take $\log v$ for integer $v>0$. For the faces $(\hat{U}_{2,5}^{1},U_{1,1}^{2})$ and $(\hat{U}_{2,5}^{1},U_{1,2}^{12})$, by restriction, they are both mapped to the face $(U_{2,3},U^{1}_{1,1})$ of $\Gamma_{3}$ on $\{2,3,4\}$, which implies that $a$ can only take $\log v$ for positive integer $v$ as well.
@@
For $F=(\hat{U}_{3,5}^{1},U_{1,1}^{1})$, $(\hat{U}_{3,5}^{1},U_{1,1}^{2})$, or $(\hat{U}_{3,5}^{1},U_{1,2}^{12})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a=\log v$ for some integer $v>0$.
@@ IV-C Entropy functions on the faces involving $U_{2,3}^{123}$
For $F=(U_{2,3}^{123},U_{1,3}^{124})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if aa= $\log{v}$ for integer $v>0$.
@@
which implies that $a$ can only take $\log{v}$ for some positive integer $k$.
The “if” part of the theorem is immediately implied by Lemma 2 and the fact that the $a=\log{v}$ in $U_{2,3}^{123}$ and whole ray $U_{1,3}^{124}$ are entropic. ∎
@@ IV-D Entropy functions on the faces involving partitions
In this subsection, we characterize the faces where the distribution of the entropy function’s characterizing random vector involves number partitions. A $t$-partition of a positive $v$ with $0\leq t\leq v-1$ is a $t$-tuple $\bm{\alpha}=(\alpha_{0},\alpha_{1},\dots,\alpha_{t-1})$ with $\alpha_{i}$ positive integer and $\sum_{i=0}^{t-1}\alpha_{i}=v$. The family of all such partitions of $v$ is denoted by $\mathcal{P}(v)$ and $H(\bm{\alpha})$ reprensent the shannon entropy $H(\frac{\alpha_{i}}{v}:i\in\mathbb{I}_{t})$.
@@
For $F=(\mathcal{W}_{2}^{12},U_{2,3}^{134})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a+b=\log{v},a=H(\dfrac{\alpha_{1}}{v},\dfrac{\alpha_{2}}{v},\dots,\dfrac{\alpha_{t}}{v})$ , where integer $v>0$ and $(\alpha_{0},\alpha_{1},\dots,\alpha_{t-1})$ is a partition of $v$.
@@
By (IV.D.27), $X_{2}$ is a function of $X_{1}$. Thus there exists a unique $x_{2}\in\mathcal{X}_{2}$ with $p(x_{1},x_{2})>0$, then $\mid\mathcal{X}_{2}\mid\leq\mid\mathcal{X}_{1}\mid=v$ and
@@ IV-E Entropy functions on the faces involving $U_{3,4}$
For $F=(U_{3,4},U^{1}_{1,1})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if aa= $\log{v}$ for integer $v>0$ .
@@
For $F=(U_{3,4},U^{12}_{1,2})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if aa= $\log{v}$ for integer $v>0$ .
@@
For $F=(U_{3,4},U^{123}_{2,3})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if a+ba+b= $\log{v}$ for integer $v>0$ .
@@
For $F=(U_{3,4},U^{123}_{1,3})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if aa= $\log{v}$ for integer $v>0$ .
@@
By Lemma 2 and the fact that $a=\log v$ on the ray $U_{3,4}$, and the whole ray $U_{1,4}$ are entropic, all $\mathbf{h}=(a,b)\in F$ are entropic when $a=\log v$ for positive integer $v$, and $b\geq 0$. In the following, we only need to show that for all $\mathbf{h}=(a,b)\in F$ are entropic, when $a>0,b>0$ with $a\neq\log v$ for some integer $v>0$.
@@ IV-F Entropy functions on the faces involving $\hat{U}_{3,5}^{1}$
For $F=(\hat{U}_{3,5}^{1},U_{1,3}^{234})$ , $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a=\log v$ for integer $v>0$.
@@
$=\sum_{x_{2},x_{3}}p(x_{2})p(x_{3})p(x_{4})=\dfrac{k_{2}}{k^{2}k_{1}},$ (IV.F.64)
where $v_{2}=\mid\{(x_{2},x_{3})|\quad p(x_{1},x_{2},x_{3},x_{4})>0\}\mid$, we conclude that $v=v_{1}v_{2}$. As $\mathbf{h}\in F$ and $(X_{i},i\in N_{4})$ is its characterizing random vector, we have
@@
For $F=(\hat{U}_{3,5}^{1},\mathcal{W}_{2}^{23})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if a+ba+b= $\log{v}$ for integer $v>0$.
@@
As for the “if” part of the theorem, we show that all the polymatroids $(a,b)$ satisfying $a+b=\log v$ are entropic. Let $p_{1},p_{2},\dots,p_{k}$ be positive value satisfying $H(p_{1},p_{2},\dots,p_{k})=a$. Let $X_{1}$ be distributed on $\mathbb{I}_{v^{2}}$ such that
$p_{X_{1}}(i)=\dfrac{p_{\lceil\frac{i+1}{k}\rceil}}{v},\text{ for }i=0,1,...,v^{2}-1.$ (IV.F.86)

### v4

@@ I Introduction
By definition, $\Gamma_{n}$ is a polyhedral cone in $\mathcal{H}_{n}$. Thus, each Shannon-type information inequality determines a face $F$ of $\Gamma_{n}$. It is natural to characterize entropy functions on the specific $F$ of $\Gamma_{n}$ (See Subsection II-B for details on the faces of a polyhedral cone). Let $F^{*}\triangleq F\cap\Gamma^{*}_{n}$ be the set of all entropy functions in $F$. In what follows, when we refer to determining the entropy functions on $F$ , or the region $F^{*}$, we will simply call this characterizing $F$. A non-Shannon-type information inequality can be considered as an outer bound on $F^{*}$ when $F=\Gamma_{n}$ itself, the improper face of $\Gamma_{n}$. A constrained non-Shannon-type information inequality is an outer bound on $F^{*}$ when $F$ is the face determined by the constraints that are equalities obtained by setting some Shannon-type inequalities as equalities. When $F$ is an extreme ray, i.e., a $1$-dimensional face of $\Gamma_{n}$, if it contains a matroid, entropy functions on $F$ are called matroidal entropy functions, and they can be fully characterized by the probabilistically characteristic set of the matroid [7, 8, 9]. Matúš fully characterized the first non-trivial 2-dimensional face of $\Gamma_{n}$ in 2006 [10]. It is a $2$-dimensional face of $\Gamma_{3}$. In 2012, Chen and Yeung characterized another $2$-dimensional face of $\Gamma_{3}$ [11]. They are the only two types of non-trivial $2$-dimensional faces of $\Gamma_{3}$ that need to be characterized. To the best of our knowledge, so far, there is no fully characterized non-trivial $3$-dimensional faces. However, partial characterizations of $3$-dimensional faces of $\Gamma_{3}$ can be found in [12, 13].
@@ II-B Faces of a polyhedral cone
$\mathbf{r}(A)=\begin{cases}\min\{2,|A|\},\quad&A\neq\{i\},\\ 2,\quad&A=\{i\},\end{cases}$
@@ II-C Entropy functions on the extreme rays of $\Gamma_{4}$
$U^{\alpha}_{2,3}$ II-C.2 [Uncaptioned image] $\{a\mathbf{r}:a=\log v,\ v\geq 1,v\in\mathbb{Z}\}$
$\mathcal{W}^{\alpha}_{2}$ II-C.2 [Uncaptioned image] $\{a\mathbf{r}:a=\log v,\ v\geq 1,v\in\mathbb{Z}\}$
$U_{2,4}$ II-C.3 [Uncaptioned image] $\{a\mathbf{r}:a=\log v,v\geq 1,v\in\mathbb{Z},v\neq 2,6\}$
$U_{3,4}$ II-C.2 [Uncaptioned image] $\{a\mathbf{r}:a=\log v,\ v\geq 1,v\in\mathbb{Z}\}$
$\hat{U}^{i}_{2,5}$ II-C.4 [Uncaptioned image] $\{a\mathbf{r}:a=\log v,\ v\geq 1,v\in\mathbb{Z}\}$
$\hat{U}^{i}_{3,5}$ II-C.5 [Uncaptioned image] $\{a\mathbf{r}:a=\log v,\ v\geq 1,v\in\mathbb{Z}\}$
@@
For the rank function $\mathbf{r}$ of $\hat{U}^{i}_{2,5}$, $\mathbf{h}=a\mathbf{r}$ is entropic if and only if $a=\log v$ for integer $v\geq 1$.
@@
For $\mathbf{h}=a\mathbf{r}$, where $\mathbf{r}$ is the rank function of $\hat{U}^{1}_{2,5}$, if we restrict it on $\{2,3,4\}$, we will see that its restriction $\mathbf{h}^{\prime}=a\mathbf{r}_{1}$, where $\mathbf{r}_{1}$ is the rank function of $U_{2,3}$, is entropic only if $a=\log v$ for integer $v\geq 1$.
@@
For the rank function $\mathbf{r}$ of $\hat{U}^{i}_{3,5}$, $\mathbf{h}=a\mathbf{r}$ is entropic if and only if $a=\log v$ for integer $v\geq 1$.
@@
By (II.C.24), (II.C.41), (II.C.34), (II.C.35), (II.C.43), (II.C.39), (II.C.40) and (II.C.42), the value of $H(X_{A})$ are determined for all $A\subseteq N$ except for $A=\{1\}$. Now we verify that $H(X_{1})=2\log v$. By (II.C.24) and (II.C.41),
@@ IV-B Entropy functions on the faces extended from $2$-dimensional faces of $\Gamma_{3}$
For $F=(U_{2,3}^{123},U_{1,1}^{1})$, $(U_{2,3}^{123},U_{1,1}^{4})$, $(U_{2,3}^{123},U_{1,2}^{14})$, $(\mathcal{W}_{2}^{14},U_{1,1}^{1})$, $(\mathcal{W}_{2}^{34},U_{1,1}^{1})$, $(\mathcal{W}_{2}^{14},U_{1,2}^{14})$ and $(\mathcal{W}_{2}^{24},U_{1,2}^{14})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a=\log{v}$ for integer $v\geq 1$.
@@
For $F=(\hat{U}_{2,5}^{1},U_{1,1}^{1})$, $(\hat{U}_{2,5}^{1},U_{1,1}^{2})$ and $(\hat{U}_{2,5}^{1},U_{1,2}^{12})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a=\log v$ for integer $v\geq 1$.
@@
By restriction, the above $3$ faces of $\Gamma_{4}$ are mapped to a subset of $\Gamma_{3}$ on $\{2,3,4\}$. The face $(\hat{U}_{2,5}^{1},U_{1,1}^{1})$ of $\Gamma_{4}$ is mapped to the extreme ray $U_{2,3}$ of $\Gamma_{3}$ on $\{2,3,4\}$, which implies that $a$ can only take $\log v$ for integer $v\geq 1$. For the faces $(\hat{U}_{2,5}^{1},U_{1,1}^{2})$ and $(\hat{U}_{2,5}^{1},U_{1,2}^{12})$, by restriction, they are both mapped to the face $(U_{2,3},U^{1}_{1,1})$ of $\Gamma_{3}$ on $\{2,3,4\}$, which implies that $a$ can only take $\log v$ for positive integer $v$ as well.
@@
For $F=(\hat{U}_{3,5}^{1},U_{1,1}^{1})$, $(\hat{U}_{3,5}^{1},U_{1,1}^{2})$, or $(\hat{U}_{3,5}^{1},U_{1,2}^{12})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a=\log v$ for some integer $v\geq 1$.
@@ IV-C Entropy functions on the faces involving $U_{2,3}^{123}$
For $F=(U_{2,3}^{123},U_{1,3}^{124})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if aa= $\log{v}$ for integer $v\geq 1$.
@@
which implies that $a$ can only take $\log{v}$ for some positive integer $v$.
The “if” part of the theorem is immediately implied by Lemma 2 and the fact that $a=\log{v}$ in $U_{2,3}^{123}$ and whole ray $U_{1,3}^{124}$ are entropic. ∎
@@ IV-D Entropy functions on the faces involving partitions
In this subsection, we characterize the faces where the distribution of the entropy function’s characterizing random vector involves number partitions. A $t$-partition of a positive $v$ with $1\leq t\leq v$ is a $t$-tuple $\bm{\alpha}=(\alpha_{0},\alpha_{1},\dots,\alpha_{t-1})$ with $\alpha_{i}$ positive integer and $\sum_{i=0}^{t-1}\alpha_{i}=v$. The family of all such partitions of $v$ is denoted by $\mathcal{P}(v)$ and $H(\bm{\alpha})$ reprensent the shannon entropy $H(\frac{\alpha_{i}}{v}:i\in\mathbb{I}_{t})$.
@@
For $F=(\mathcal{W}_{2}^{12},U_{2,3}^{134})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a+b=\log{v},a=H(\dfrac{\alpha_{0}}{v},\dfrac{\alpha_{1}}{v},\dots,\dfrac{\alpha_{t-1}}{v})$ , where integer $v\geq 1$ and $(\alpha_{0},\alpha_{1},\dots,\alpha_{t-1})$ is a partition of $v$.
@@
By (IV.D.27), $X_{2}$ is a function of $X_{1}$. Thus there exists a unique $x_{2}\in\mathcal{X}_{2}$ with $p(x_{1},x_{2})>0$, then $|\mathcal{X}_{2}|\leq|\mathcal{X}_{1}|=v$ and
@@ IV-E Entropy functions on the faces involving $U_{3,4}$
For $F=(U_{3,4},U^{1}_{1,1})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if aa= $\log{v}$ for integer $v\geq 1$ .
@@
For $F=(U_{3,4},U^{12}_{1,2})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if aa= $\log{v}$ for integer $v\geq 1$ .
@@
For $F=(U_{3,4},U^{123}_{2,3})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if a+ba+b= $\log{v}$ for integer $v\geq 1$ .
@@
For $F=(U_{3,4},U^{123}_{1,3})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if aa= $\log{v}$ for integer $v\geq 1$ .
@@
By Lemma 2 and the fact that $a=\log v$ on the ray $U_{3,4}$, and the whole ray $U_{1,4}$ are entropic, all $\mathbf{h}=(a,b)\in F$ are entropic when $a=\log v$ for positive integer $v$, and $b\geq 0$. In the following, we only need to show that for all $\mathbf{h}=(a,b)\in F$ are entropic, when $a>0,b>0$ with $a\neq\log v$ for some integer $v\geq 1$.
@@ IV-F Entropy functions on the faces involving $\hat{U}_{3,5}^{1}$
For $F=(\hat{U}_{3,5}^{1},U_{1,3}^{234})$ , $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a=\log v$ for integer $v\geq 1$.
@@
$=\sum_{x_{2},x_{3}}p(x_{2})p(x_{3})p(x_{4})=\dfrac{v_{2}}{v^{2}v_{1}},$ (IV.F.64)
where $v_{2}=|\{(x_{2},x_{3})|\quad p(x_{1},x_{2},x_{3},x_{4})>0\}|$, we conclude that $v=v_{1}v_{2}$. As $\mathbf{h}\in F$ and $(X_{i},i\in N_{4})$ is its characterizing random vector, we have
@@
For $F=(\hat{U}_{3,5}^{1},\mathcal{W}_{2}^{23})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if a+ba+b= $\log{v}$ for integer $v\geq 1$.
@@
As for the “if” part of the theorem, we show that all the polymatroids $(a,b)$ satisfying $a+b=\log v$ are entropic. Let $p_{1},p_{2},\dots,p_{v}$ be positive value satisfying $H(p_{1},p_{2},\dots,p_{v})=a$. Let $X_{1}$ be distributed on $\mathbb{I}_{v^{2}}$ such that
$p_{X_{1}}(i)=\dfrac{p_{\lceil\frac{i+1}{v}\rceil}}{v},\text{ for }i=0,1,...,v^{2}-1.$ (IV.F.86)